OpenStudy (anonymous):

56 krissywatts I need math help a.s.a.p on perimeter of a rectangle.? the perimeter of a rectangle is 46 feet. The width is 3ft less than twice the length. Find the length and width of the rectangle. step by step is appreciated so i can see where i went wrong.

OpenStudy (agent47):

P=46 w=2l-3

OpenStudy (agent47):

2(w+l)=46 2(2l-3+l)=46 3l-3=23 3l=26

OpenStudy (agent47):

l=26/3 w = 2*26/3 - 3

OpenStudy (anonymous):

its not coming out right

OpenStudy (anonymous):

P=2(L)+2(w) P=46 w=2x-3 L=2x 46=2(2x)+2(2x-3) 46=4x+4x-6 8x=52 x=52/8=13/2=6 1/2=6.5 L=2*6.5=13 w=2x-3=13-3=10 13 x 10 thats a 13 by 10

OpenStudy (anonymous):

thanks